This question is a bit confusing, can anyone help explain?

Question: "A well-known piano manufacturer wishes to expand in China. It decides that 80% of the $32 million it needs will come from debt, and the remaining 20% from selling equity. The cost of debt is 7% and the corporate tax is 35%. To estimate the cost of equity, the firm uses the CAPM with these parameters: equation

o What is the WACC, or hurdle rate, of the investment?

o If the IRR of the investment is estimated at 7 %, should it proceed?
Answer #1

Here,

Manufacturer of piano wants to expand in China, and for this he will need $32 millions

He wants to source 80% debt and 20% equity,

So debt required = 0.8*32 = $25.6 million

Equity sourced = $6.4 million

Cost of debt = 7%

Tax rate = 35%

To estimate cost of equity firm uses CAPM,

re = Rf + Beta(Rm - Rf)

WACC = wd *cd(1-t) + we*ce

If WACC is smaller than IRR of 7% then only firm should for the project.